Find the indicated IQ score. The graph to the right depicts IQ scores of​ adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15. with the distribution of 0.75


1
Expert's answer
2022-02-28T09:12:41-0500
"P(X>x)=1-P(Z\le\dfrac{x-\mu}{\sigma})"

"=1-P(Z\le\dfrac{x-100}{15})=0.75"

"P(Z\le\dfrac{x-100}{15})=0.25"

"\dfrac{x-100}{15}\approx-0.6745"

"x=89.88"